# femtoPro: Real-time linear and nonlinear optics simulations

### v1

@@ 3.2 N-beam interference
We proceed in the coordinate system in which the $\hat{\mathbf{z}}$ axis is parallel to the average of the incident wave vectors. As will be derived below, interference is visible only for small phase mismatch, and thus the approximation is valid that the beam positions, $(x_{k},y_{k})$ and $(x_{l},y_{l})$, and radii, $w_{k}$ and $w_{l}$, need not be transformed because the average direction is almost identical to either of the incident directions. The actual incidence angle onto the detector plane is not decisive in first order because the underlying spatial interference pattern is retained, and thus integration over the tilted pattern provides a comparable total power as integration over the normal-incidence pattern.
@@
$\begin{split}\eta_{k,l}=&\frac{2w_{k}w_{l}}{w_{k}^{2}+w_{l}^{2}}\exp\left[-\frac{(x_{k}-x_{l})^{2}+(y_{k}-y_{l})^{2}}{w_{k}^{2}+w_{l}^{2}}\right]\\ &\times e^{i(\Delta k_{x}x_{\text{p}}+\Delta k_{y}y_{\text{p}})}\exp\left[-\frac{\Delta k_{x}^{2}+\Delta k_{y}^{2}}{2\left(\frac{1}{w_{k}^{2}}+\frac{1}{w_{l}^{2}}\right)}\right].\end{split}$ (28)
@@
In Fig. 4(a), we illustrate the dependence of the interference visibility $\eta_{1,2}$ between two beams on various parameters. In the first example [Fig. 4(a)], one beam remains constant and the second beam is shifted laterally with otherwise identical parameters. We show three exemplary curves for the cases in which both beams have a radius of 1 mm (blue), 2 mm (red), and 4 mm (green). It is evident that the interference visibility can be maintained over a larger lateral shift range in case of larger beam radii. In Fig. 4(b), we show the influence of the wave-vector mismatch on the interference visibility by varying the propagation angle of the second beam relative to the first. Here, the angular mismatch has less impact on the interference visibility when both beams have smaller radii, which reflects the general behavior of the last exponential function in Eq. (28).
Figure 4: Interference visibility factor $\eta_{1,2}$ between two light fields. (a) Interference visibility resulting from the second beam being laterally shifted relative to the first beam where both beams have a radius of 1 mm (blue), 2 mm (red) or 4 mm (green). (b) Dependence of the interference visibility on the relative incidence angle of the second beam [colors correspond to the same beam radii as in (a)]. (c) Interference visibility of two beams incident on an aperture with varying radius. Here, both beams have a radius of 4 mm and one of the beams is laterally shifted by 1 mm from the center of the aperture at normal incidence while the other beam is laterally shifted by 4 mm relative to the first beam and has an incidence angle of 0.01 (fixed parameters).
In Section LABEL:SISecInterferenceVisibilityFactor of the SI, we discuss quantitatively how interference visibility can be increased by closing an entrance aperture $a$ directly in front of a spectrometer, and we find the conditions for maximal angular mismatch and transverse misalignment between incident beams that still ensure interference. In Fig. 4(c), we display the interference visibility factor in the case where two beams with a radius of 4 mm are incident on an aperture while the size of the aperture is changed. One of the beams is laterally shifted by 1 mm relative to the center of the aperture at normal incidence whereas the other beam has an exemplary lateral shift of 4 mm relative to the first beam and is incident at an angle of $0.01^{\circ}$. The two beams hit the aperture at a distance of 1 m from their origin, resulting in the second beam being displaced in the plane of the iris by 4.8 mm relative to the center of the iris. This configuration leads to a beam overlap region that is displaced relative to the center of the iris, as indicated in the inset of Fig. 4(c). This example shows that closing the aperture (initially considered to have an infinite opening radius) leads to an increase of the interferometric contrast as those parts of the beams that do not overlap get truncated. However, when closing the aperture further, the overlapping parts get truncated as well, which leads to reduction in contrast such that there is an optimum aperture opening. Note that the slight kink at an aperture radius of 3 mm is due to the transition in how the beam cutting is calculated as discussed in connection with Fig. 3(b).
@@ 4 Modulations from light–matter interaction
4.1 First-order non-resonant response
@@
and apply only the remaining terms $b_{2,\text{disp}}$ and $b_{3,\text{disp}}$ in the non-resonant part of the linear modulation function, $M_{\text{non-resonant}}(\omega)$. Note that concerning the geometry, we work in the limit of infinitesimally thin GOEs with one central “active plane,” so that the incident and outgoing beams have already been assumed to travel in vacuum up to and starting from that plane. Thus, we need to subtract $L/c$ from the (modified) propagation time because it has already been accounted for, leading to the term “$-1$” in Eq. (31). In the case of reflection, finite values for the group-delay dispersion or the third-order phase coefficient may be provided directly to describe, e.g., chirped mirrors.
4.2 First-order resonant response
It is convenient to start with a response function in time domain, rather than frequency domain, to treat a resonantly excited material such as an ensemble of molecules in a solvent. The frequency-domain modulation function is then obtained by Fourier transformation via Eq. (LABEL:SIEqEomegaJFourier) and appropriate further scaling, as described below.
The linear temporal response function in a Franck–Condon model under the rotating-wave approximation (RWA) is given by [24]
$S_{\text{RWA}}^{(1)}(t)=\theta(t)\frac{i}{\hbar}\mu_{\text{eg}}^{2}\exp\left(-\gamma t-\frac{1}{2}\Delta\omega^{2}t^{2}\right)e^{-i(\omega_{\text{eg}}-\omega_{0})t}e^{-S_{\text{HR}}}\sum_{j=0}^{N_{\text{vib}}-1}\frac{S_{\text{HR}}^{j}}{j!}e^{-ij\omega_{\text{vib}}t},$ (32)
[…]
We now have to introduce an appropriate scaling of $\mu_{\text{eg}}$ to convert between “physical” quantities and computationally stored arrays. This problem is similar to the scaling of electric fields that we solved by normalizing with respect to a given temporal power and, ultimately, a given pulse energy, via Eq. (LABEL:SIEqTemporalPowerDefinition). Thus we apply a similar strategy here. Furthermore, we seek a behavior of the modulation function such that longer paths through the material lead to stronger absorption fulfilling the Lambert–Beer law for the pulse energy,
$W_{\text{out}}=W_{\text{in}}e^{-\alpha L},$ (33)
with a given extinction coefficient $\alpha$ of the material.
As a first step, we evaluate Eq. (32) on the same time grid as for regular laser pulses [Eq. (LABEL:SIEqt(j))] and set $\mu_{\text{eg}}^{2}/\hbar=1$. The Fourier transform of the result is the (unscaled) frequency-domain linear susceptibility,
$\chi_{\text{unscaled}}(\omega)=\operatorname{\mathfrak{F}}\left[\left.S_{\text{RWA}}^{(1)}(t)\right|_{\mu_{\text{eg}}^{2}/\hbar=1}\right],$ (34)
from which we get the (unscaled) complex index of refraction,
$\tilde{n}_{\text{unscaled}}(\omega)=\sqrt{\chi_{\text{unscaled}}(\omega)+1},$ (35)
where it is understood that the square-root symbol indicates a complex-valued square-root operation. Now we obtain a suitable scaling factor,
$s=\frac{\alpha c}{2\omega_{\alpha}\operatorname{Im}\tilde{n}_{\text{unscaled}}(\omega_{\alpha})},$ (36)
where $\alpha$ is the absorption coefficient of Eq. (33) to which we want to scale the material response and $\omega_{\alpha}$ is the angular frequency at which this absorption coefficient shall be reached (typically a peak of the resulting absorption spectrum that can be found by locating the frequency at which $\operatorname{Im}\tilde{n}_{\text{unscaled}}(\omega)$ has a maximum). Using Eq. (36) allows us to calculate a scaled susceptibility,
$\chi(\omega)=s\chi_{\text{unscaled}}(\omega),$ (37)
and a scaled complex index of refraction,
$\tilde{n}(\omega)=\sqrt{\chi(\omega)+1},$ (38)
and finally the desired modulation function,
$M_{\text{resonant}}(\omega)=\exp\left(i\frac{\tilde{n}(\omega)\omega L}{c}\right),$ (39)
for a material of thickness $L$.
[…]
Figure 5: Linear resonant response. Exemplary modulation function $M_{\text{resonant}}(\omega)$ for an absorbing molecular response (red: real part, blue: imaginary part).
@@ 4.3 Second-order response
where “c.c.” indicates the complex conjugate of the previous term. This is the level of treatment often found in didactic textbooks on second-order response. However, we need to consider also the spatial properties of the beams if we want to simulate and understand the effect of beam alignment on nonlinear phenomena. Here we derive the explicit equations in terms of the computationally stored complex envelope arrays $\tilde{E}_{t}(j)$ and associated parameters, we take into account the spatial beam profile, and we determine the missing proportionality factor between $E_{\text{s}}(t)$ and $E^{2}(t)$.
Equation (40) contains multiplicative electric-field terms. Hence, we use Eq. (1) and multiply the propagated fields,
$\begin{split}&mE_{1,\text{prop}}^{+}(x,y,z,t)E_{2,\text{prop}}^{+}(x,y,z,t)\\ =\,&m\sqrt{\frac{2}{\pi}}\frac{1}{w_{1}(z)}\exp\left[-\frac{(x-x_{1})^{2}+(y-y_{1})^{2}}{w_{1}^{2}(z)}\right]e^{i\mathbf{k}_{1}\cdot\mathbf{r}}\sqrt{\frac{S_{1}}{\delta t}}\tilde{E}_{1,t}(j_{1})e^{-i\omega_{1}(t-T_{1})}\\ &\times\sqrt{\frac{2}{\pi}}\frac{1}{w_{2}(z)}\exp\left[-\frac{(x-x_{2})^{2}+(y-y_{2})^{2}}{w_{2}^{2}(z)}\right]e^{i\mathbf{k}_{2}\cdot\mathbf{r}}\sqrt{\frac{S_{2}}{\delta t}}\tilde{E}_{2,t}(j_{2})e^{-i\omega_{2}(t-T_{2})},\end{split}$ (41)
wherein the multiplication of the two Gaussian exponentials generates a new Gaussian $A_{\text{p}}(x,y)$ according to Eq. (LABEL:SIEqGaussianProduct) of the SI with the product amplitude $A_{\text{p}}$ in Eq. (LABEL:SIEqProductAmplitude) and $A_{1}=A_{2}=1$. Because we consider the stigmatic approximation, we have $w_{1,x}=w_{1,y}$ and $w_{2,x}=w_{2,y}$ according to Eq. (LABEL:SIEqGaussianProduct) to Eq. (LABEL:SIEqProductAmplitude) of the SI. Further, $m=1$ for second-harmonic generation (SHG) but $m=2$ for sum-frequency generation (SFG) according to Eq. (40) due to the two possible permutations $E_{1}E_{2}$ and $E_{2}E_{1}$ that both deliver an SFG field in the same direction.
We want to write the result of Eq. (41) as a new (but not yet correctly energy-scaled) SFG field,
$\begin{split}E_{\text{SFG,prop}}^{+}(x,y,z,t)=\,&\sqrt{\frac{2}{\pi}}\frac{1}{w_{\text{SFG}}}\exp\left[-\frac{(x-x_{\text{SFG}})^{2}+(y-y_{\text{SFG}})^{2}}{w_{\text{SFG}}^{2}}\right]e^{i\mathbf{k}_{\text{SFG}}\cdot\mathbf{r}}\\ &\times\sqrt{\frac{S_{\text{SFG,unscaled}}}{\delta t}}\tilde{E}_{\text{SFG},t}(j_{\text{SFG}})e^{-i\omega_{\text{SFG}}(t-T_{\text{SFG}})},\end{split}$ (42)
with appropriate parameters. Comparing the individual factors of Eq. (41) and Eq. (42), we find:
[…]
$\omega_{\text{SFG}}$ $=\omega_{1}+\omega_{2},$ (46)
$\mathbf{k}_{\text{SFG}}$ $=\mathbf{k}_{1}+\mathbf{k}_{2},$ (47)
$T_{\text{SFG}}$ $=\frac{\omega_{1}T_{1}+\omega_{2}T_{2}}{\omega_{1}+\omega_{2}},$ (48)
$\tilde{E}_{\text{SFG},t}(j_{\text{SFG}})$ $=\tilde{E}_{1,t}(j_{1})\tilde{E}_{2,t}(j_{2}),$ (49)
$\sqrt{\frac{2}{\pi}}\frac{1}{w_{\text{SFG}}}\sqrt{\frac{S_{\text{SFG,unscaled}}}{\delta t}}$ $=m\frac{2}{\pi}\frac{1}{w_{1}(z)w_{2}(z)}A_{\text{p}}\frac{\sqrt{S_{1}S_{2}}}{\delta t}.$ (50)
Note that the beam radius of the SFG beam is smaller than those of the fundamental beams, as expected due to the field multiplication. If both incident beams have the same radius, they are larger than the SFG radius by a factor of $\sqrt{2}$ according to Eq. (45). Solving Eq. (50) for $S_{\text{SFG,unscaled}}$, using the explicit results for $w_{\text{SFG}}$ from Eq. (45) and for $A_{\text{p}}$ from Eq. (LABEL:SIEqProductAmplitude) (with $A_{1}=A_{2}=1$) and simplifying leads to
$S_{\text{SFG,unscaled}}=\frac{2m^{2}}{\pi}\frac{1}{w_{1}^{2}(z)+w_{2}^{2}(z)}\frac{S_{1}S_{2}}{\delta t}\exp\left[-2\frac{(x_{1}-x_{2})^{2}+(y_{1}-y_{2})^{2}}{w_{1}^{2}(z)+w_{2}^{2}(z)}\right].$ (51)
We illustrate how $S_{\text{SFG,unscaled}}$ varies when displacing one of the fundamental beams relative to the other in Fig. 6. Each of the fundamental beams has a pulse energy of 1 mJ and the beam radius of one beam was set to 1 mm (blue), 2 mm (red), and 4 mm (green) while that of the other beam is kept at 1 mm. It is evident that the scaling factor generally shows the largest change with increasing overlap when both beams have small radii. In other words, the SFG intensity at maximum overlap increases with increasing energy density of the fundamental beams.
Figure 6: Beam energy scaling factor for sum-frequency generation. The plots show the dependence of the beam energy scaling factor for sum-frequency generation, $S_{\text{SFG,unscaled}}$, on the lateral shift between two fundamental beams. The beam radius of the first beam was varied (blue: 1 mm, red: 2 mm, green: 4 mm), whereas that of the second beam was held constant (1 mm).
Concerning the multiplication of envelopes in Eq. (49), we have to use the appropriate array indices. Note that we have to add the respective propagation times $T_{1}$ and $T_{2}$ to Eq. (LABEL:SIEqt(j)) before extracting the relevant indices $j_{1}$ and $j_{2}$ of the array elements to be multiplied in Eq. (49). In the multiplication, it will thus happen that a theoretically requested array index is out of range of the available indices $j=\{0,1,\ldots,N_{\text{s}}-1\}$, i.e., the time shift between the two pulses is such that for a time coordinate $t(j_{1})$ at index $j_{1}$ of the first field no “matching” index $j_{2}$ is found because the particular $t(j_{1})$ is not contained in the list for $t(j_{2})$ of the second field. Then we assume the multiplier field to be zero, and thus also the SHG field will vanish at this time coordinate.
Now we have to find an appropriate scaling factor for the SFG field. In principle, proper perturbation theory of light–matter interaction delivers already the correct scaling factor of the absolute signal strength in terms of an appropriate response function. This is, however, not applicable in our situation because we chose a simplified treatment for reasons of computational speed, rather than dealing with a more realistic but much more complicated full spatial–temporal model [40, 41, 42, 43, 44]. In particular, the response-function treatment provides the field at one particular spatial position only, and one then has to solve the wave equation with that nonlinear field as a source term upon propagation through the full length of the material. This is beyond our scope. Thus, we do not take into account propagation effects and want the final signal to scale simply with the square of the incident field. Then we have to ensure, however, that the pulse energy in the second-harmonic field does not exceed the sum of pulse energies in the initial fields because then energy conservation could not be fulfilled. In practice, we will require only a (small) fraction of the power to be converted in order to prevent saturation effects. We derive in Section LABEL:SISecAbsoluteAmplitudeScalingOfSecondOrderFields of the SI an appropriate scaling factor.
If we allow only a fraction $\eta_{2}$ of the fundamental pulse energy $W_{0}$ to be converted to SHG under optimum conditions (leading to a maximum SHG pulse energy $W_{\text{SHG,max}}$), the final result for the pulse-energy scaling factor is
$S_{\text{SFG}}=\eta_{2}\frac{W_{0}}{W_{\text{SHG,max}}}S_{\text{SFG,unscaled}}.$ (52)
We may set $0<\eta_{2}\leq 1$ at will to guarantee energy conservation in conjunction with an appropriate reduction of the fundamental energies as explained in Section 4.5. We may select, e.g., $\eta_{2}=10\,\mathrm{\%}$ to stay away from the saturation regime. This evaluation has to be done only once for a given laser (unless the laser output parameters are changed). The intermediate multiplication result $\eta_{2}W_{0}/W_{\text{SHG,max}}$ can then simply be stored as a number and used as a multiplier on $S_{\text{SHG,unscaled}}$.
4.4 Nonlinear signal beam curvature
We now discuss the beam radius and radius of curvature of a nonlinear signal beam that arises from the corresponding properties of the incident fields. The incident fields may be mutually different in general, e.g., in pump–probe experiments with different beam radii of pump and probe beams. We thus need to address where the focus of the resulting beam will be located. A detailed derivation is presented in Section LABEL:SISecDerivationOfNonlinearSignalBeamCurvature of the SI, and we only quote the results here.
It turns out that the $n$th-order nonlinear signal beam radius, $w_{\text{s}}(z)$, at the position of a thin sample, $z$, is obtained from the individual beam radii of the $n$ incident beams, $w_{i}(z)$, via
$\frac{1}{w_{\text{s}}^{2}(z)}=\sum_{i=1}^{n}\frac{1}{w_{i}^{2}(z)},$ (53)
in generalization of Eq. (45). The curvature radius of the signal beam at the sample position, $R_{\text{s}}(z)$, is given by
$R_{\text{s}}(z)=\frac{\pi w_{\text{s}}^{2}(z)}{\lambda\zeta_{\text{s}}}$ (54)
with the longitudinal position parameter
$\zeta_{\text{s}}=w_{\text{s}}^{2}(z)\sum_{i=1}^{n}\frac{\alpha_{i}\zeta_{i}}{w_{i}^{2}(z)}$ (55)
with phase-matching coefficients $\alpha_{i}\in\{-1,+1\}$ of the incident beams indicating with which sign they contribute to the calculation of the nonlinear signal phase-matching direction, and
$\zeta_{i}=-\frac{\operatorname{Re}q_{i}}{\operatorname{Im}q_{i}}$ (56)
using the complex beam parameters $q_{i}$ of the incident beams [compare Eq. (LABEL:SIEqComplexRadiusOfCurvature)].
Figure 7: Nonlinear signal beam. (a) Lateral position of the SHG beam waist emerging from two collinear fundamental beams with the same radius while the second beam is laterally shifted. The lateral position of the SHG beam is given relative to that of the first fundamental beam. (b) Beam waist of the SHG beam depending on the beam waist of the second fundamental beam. The waist of the first fundamental beam is set to 1 mm.
For one particular limiting case, consider that the sample or nonlinear crystal is placed near the position of the beam waist of all incident beams, i.e., $\zeta_{i}\ll 1$. This represents the “plane-wave” situation (for which $R_{i}\to\infty$) that can be realized either by placing the sample at the focal point or by employing collimated beams. Then, $\zeta_{\text{s}}\ll 1$, and the signal focus is located at the same longitudinal position as the incident beams are. In other words, the signal beam is created at its waist in the sample.
In Fig. 7(a), we illustrate different focusing conditions of two incoming beams by plotting the lateral SHG beam waist position while varying the lateral displacement between the incoming beams. We consider two collinear beams under perfect phase matching. In case of perfect co-focusing, the waist position of the SHG beam is the same as that of the incident beams. When one of the incident beams is laterally displaced, the SHG beam waist position also shifts, but by half the value of the displacement. For SHG, the Rayleigh range $z_{\text{R}}$ is the same as that of the fundamental [22]. Thus, the half-divergence angle of SHG is smaller by $1/\sqrt{2}$ with respect to the fundamental according to Eq. (LABEL:SIEqHalfDivergenceAngle), with the beam waist smaller by the same factor. If the two beams are co-focused on the sample while the first beam has a fixed radius and the radius of the other beam is increased, the radius of the SHG beam approaches that of the first beam asymptotically [Fig. 7(b)].
Consider for further illustration, as another exemplary special case, pump–probe transient absorption spectroscopy, even though its response-function treatment is beyond the scope of the present work. Nevertheless, this case is instructive for illuminating the generality of the derived beam-parameter results. In that case, the phase-matching direction of the signal is given by
$\mathbf{k}_{\text{s}}=-\mathbf{k}_{\text{pu}}+\mathbf{k}_{\text{pu}}+\mathbf{k}_{\text{pr}}$ (57)
for incident pump ($\mathbf{k}_{\text{pu}}$) and probe ($\mathbf{k}_{\text{pr}}$) wave vectors. The wave vector $\mathbf{k}_{\text{pu}}$ appears twice in this equation, once with a positive sign and once with a negative sign, because there are two (phase-conjugate) interactions with the pump pulse. Thus, the phase-matching coefficients are $\{\alpha_{1},\alpha_{2},\alpha_{3}\}=\{-1,+1,+1\}$ and the signal propagates into the direction of the probe beam (leading to constructive or destructive interference that is then interpreted in measurements as absorbance changes). From Eq. (53), we obtain
$\frac{1}{w_{\text{s}}^{2}(z)}=\frac{2}{w_{\text{pu}}^{2}(z)}+\frac{1}{w_{\text{pr}}^{2}(z)}.$ (58)
It is often customary to choose the pump-beam radius (much) larger than the probe-beam radius. Under this condition, we find from Eq. (58) that $w_{\text{s}}\approx w_{\text{pr}}$ and from Eq. (55) that $\zeta_{\text{s}}\approx\zeta_{\text{pr}}$, and thus the signal beam has the exact same characteristics as the probe beam. This is the reason for choosing a larger pump-beam diameter because then the interference leads to a spatially uniform transient absorption signature across the full probe-beam profile. Otherwise, the signal beam might deviate from the probe beam and one loses contrast.
4.5 Reduction of fundamental pulse energies upon nonlinear signal generation
@@
When new signal fields are generated according to second-order or third-order response and no energy is absorbed in the nonlinear crystal or sample, the total energy summed up over all laser beams should be conserved. In a fully self-consistent solution of Maxwell’s equations, this would arise automatically. Our simplified treatment calls for manual adjustment of fundamental pulse energies. Otherwise, it might happen that the total energy of all output beams after a nonlinear GOE exceeds the total energy of all incident beams.
We arrive at the correct energy, contributed to the new beam by each fundamental, by analyzing the relevant number of photons for each microscopic signal-generation process. Let us consider sum-frequency generation (SFG) as in Eq. (40). Each SFG photon has an energy of $\hbar\omega_{\text{SFG}}$, and the fundamental beams $\tilde{E}_{1}(t)$ and $\tilde{E}_{2}(t)$ carry photons of energy $\hbar\omega_{1}$ and $\hbar\omega_{2}$, respectively. The incident fundamental pulse energies, $W_{1,\text{in}}$ and $W_{2,\text{in}}$, and the generated SFG pulse energy, $W_{\text{SFG}}$, can be obtained from the fields via Eq. (LABEL:SIEqEnergyIsScaleTimesNormSquared) from the SI. Thus, a number of $W_{\text{SFG}}/(\hbar\omega_{\text{SFG}})$ photons have to be generated. This requires the same number of photons contributed from each of the incident beams, reducing their energies accordingly by
$\Delta W_{1}=\frac{\hbar\omega_{1}}{\hbar\omega_{\text{SFG}}}W_{\text{SFG}},\quad\Delta W_{2}=\frac{\hbar\omega_{2}}{\hbar\omega_{\text{SFG}}}W_{\text{SFG}}.$ (59)
Analogously, one obtains the beam reductions for second-harmonic generation by setting $W_{2}=W_{1}$, i.e., the fundamental beam is reduced by twice the value (in addition to the reduction from potential SFG contributions). Furthermore, it is possible that more than two beams interact in any given GOE with second-order response. In that case, we calculate the reductions for all beam-pair combinations (SFG and SHG) and then, in the end, obtain the reduced fundamental output pulse energies by
$W_{1,\text{out}}=W_{1,\text{in}}-\sum_{k}\Delta W_{1,k},\quad W_{2,\text{out}}=W_{2,\text{in}}-\sum_{k}\Delta W_{2,k},$ (60)
where the index $k$ runs over all pairwise combined processes. Note that we reduce the fundamental pulse energies only after having generated all second-order beams with the full incident energy. This ensures that no particular pair of beams is “prioritized” and has a higher pulse energy than any of the others when calculating the SFG and SHG beams successively. Of course the beam-energy reductions occur in addition to those from beam clipping as discussed in Section 3.1, and we do not modify $W$ directly but rather the pulse-energy scaling factor $S$ from Eq. (LABEL:SIEqEnergyIsScaleTimesNormSquared). Working with photon numbers and center frequencies only is an approximation. In principle, one could analyze the complete spectral intensity shape for the correct result [32], but since we require only the adjustment of absolute energies here, such small variations are irrelevant. For the spectral–temporal shapes of the signal and transmitted pulses themselves we take into account the explicit field profiles.

### v2

@@ 3 N-beam interference
We proceed in the coordinate system in which the $\hat{\mathbf{z}}$ axis is parallel to the average of the incident wave vectors. As will be derived in the SI, interference is visible only for small phase mismatch, and thus the approximation is valid that the beam positions, $(x_{k},y_{k})$ and $(x_{l},y_{l})$, and radii, $w_{k}$ and $w_{l}$, need not be transformed because the average direction is almost identical to either of the incident directions. The actual incidence angle onto the detector plane is not decisive in first order because the underlying spatial interference pattern is retained, and thus integration over the tilted pattern provides a comparable total power as integration over the normal-incidence pattern.
@@
$\eta_{k,l}=\frac{2w_{k}w_{l}}{w_{k}^{2}+w_{l}^{2}}\exp\left[-\frac{(x_{k}-x_{l})^{2}+(y_{k}-y_{l})^{2}}{w_{k}^{2}+w_{l}^{2}}\right]e^{i(\Delta k_{x}x_{\text{p}}+\Delta k_{y}y_{\text{p}})}\exp\left[-\frac{\Delta k_{x}^{2}+\Delta k_{y}^{2}}{2\left(\frac{1}{w_{k}^{2}}+\frac{1}{w_{l}^{2}}\right)}\right].$ (10)
@@ 4 Modulations from light–matter interaction
4.1 First-order response
@@
and apply only the remaining terms $b_{2,\text{disp}}$ and $b_{3,\text{disp}}$ in the non-resonant part of the linear modulation function, $M_{\text{non-resonant}}(\omega)$. Note that concerning the geometry, we work in the limit of infinitesimally thin GOEs with one central “active plane,” so that the incident and outgoing beams have already been assumed to travel in vacuum up to and starting from that plane. Thus, we need to subtract $L/c$ from the (modified) propagation time because it has already been accounted for, leading to the term “$-1$” in Eq. (13). In the case of reflection, group-delay dispersion or a third-order phase coefficient may be provided directly to describe, e.g., chirped mirrors.
We treat the case of first-order resonant response, on the example of a molecular sample with an electronic transition coupled to vibrational modes, in Section LABEL:SISecFirstOrderResonantResponse of the SI.
@@ 4.2 Second-order response
where “c.c.” indicates the complex conjugate of the previous term. This is the level of treatment often found in didactic textbooks on second-order response. However, we need to consider also the spatial properties of the beams if we want to simulate and understand the effect of beam alignment on nonlinear phenomena. In Section LABEL:SISecDerivationSecondOrderResponse of the SI, we derive the explicit equations in terms of the computationally stored complex envelope arrays $\tilde{E}_{t}(j)$ and associated parameters, we take into account the spatial beam profile, and we determine the missing proportionality factor between $E_{\text{s}}(t)$ and $E^{2}(t)$.
The correct description of the nonlinear signal field further requires to calculate its beam radius and the radius of curvature, depending on the incident fields. These relations are derived in Section LABEL:SISecNonlinearSignalBeamCurvature of the SI.
@@
When new signal fields are generated according to second-order or third-order response and no energy is absorbed in the nonlinear crystal or sample, the total energy summed up over all laser beams should be conserved. In a fully self-consistent solution of Maxwell’s equations, this would arise automatically. Our simplified treatment calls for manual adjustment of fundamental pulse energies. Otherwise, it might happen that the total energy of all output beams after a nonlinear GOE exceeds the total energy of all incident beams. We arrive at the correct energy, contributed to the new beam by each fundamental, by analyzing the relevant number of photons for each microscopic signal-generation process as derived in Section LABEL:SISecReductionOfFundamentalPulseEnergiesUponNonlinearSignalGeneration of the SI.